An ideal gas thermometer and a mercury thermometer are calibrated at 0 oC and at 100 oC. The thermal expansion coefficient for mercury is 𝛼 = 1 𝑉0 ( πœ•π‘‰ πœ•π‘‡ ) 𝑃 = 1.817π‘₯10βˆ’4 + 5.90π‘₯10βˆ’9πœƒ + 3.45π‘₯10βˆ’10πœƒ2 , where πœƒ is the value of the Celsius temperature and 𝑉0 = 𝑉 at πœƒ = 0. What temperature would appear on the mercury scale when the ideal gas scale reads 50 oC?Β 


1
Expert's answer
2021-05-09T15:12:33-0400

Given,

"T_i=0^\circ C"

"T_f=100 ^\circ C"

"\alpha = \frac{1}{V_o}(\frac{\partial V}{\partial T})_P =1.817\times 10^{-4}+5.90 \times 10^{-9}\theta +3.45\times 10^{-10}\times{\theta^2}"


"\Rightarrow \alpha = \frac{1}{V_o}(\frac{\partial V}{\partial T})_P"

"\Rightarrow \frac{\partial V}{V_o}=\alpha \partial T"

Now, taking the integration of both side,

"\int\frac{\partial V}{V_o}=\int(1.817\times 10^{-4}+5.90 \times 10^{-9}\theta +3.45\times 10^{-10}\times{\theta^2}){\partial T}"

"\Rightarrow \ln(V_o)=1.817\times 10^{-4}T+\frac{5.90\times 10^{-9} T^2}{2}+3.45\times 10^{-10}\frac{T^3}{3}"